# Anatomy of Elite and Mass Polarization in Social Networks

### v1

@@ 3.3 Elite is significantly more aligned than the mass
Elite members are consistently more aligned than the public in all topics studied here, as shown in Figure 5A. Higher alignment suggests that the elite regularly maintain more predictable stances on various political themes, as typically expected. In 2019, among both elite and mass, topic pairs immigration & education and immigration & climate were the most aligned. Conversely, the lowest alignment was found between the topics immigration & social and education & social.
@@
After four years the gap between elite and mass alignment widened – not only did the elite show a much higher level of issue alignment (the mean level doubled), but the disparity between the mass and the elite also increased in 2023. The overall level of mass alignment did not stay at the same level as in 2019 either. Instead, the mass also experienced a notable increase in alignment across all topics, as the mean line, corresponding to the average of alignment over all the topics, rose from 0.13 to 0.44. For the elite groups, eight out of ten pairs of topics had an NMI value over 0.8, which can be considered very high. Results show that knowing elite member’s stance on climate politics would reveal the most information on their views about immigration policies compared to the other topics included in this study. Interestingly, climate & economy experienced the smallest increase in alignment.
To evaluate deeper the relationship between the elite and mass alignment, we employ bayesian linear regression. As our study contains five distinct topics, the total number of pairs are ten for each year, resulting in a total of twenty samples for the entire period. We fit a linear regression using elite alignment value as an independent variable and mass alignment as dependent variable (see Figure 6). The bayesian framework enables the quantification of uncertainty in model parameters, which we find crucial due to the small sample size. While we acknowledge that our approach may oversimplify the relationship between the elite and mass, we believe that our analysis sheds light on the behavior of these distinct hierarchical groups and highlights how the interdependence between the variables has intensified over time. To be more specific, here we consider the slope of the regression to correspond to the strength of the relationship between the elite and mass alignment. It’s also worth reminding that the linear dependence between the variables we are about to report doesn’t necessarily imply causality. However, the slope of linear regression is tightly connected to the pearson correlation coefficient, as
$\text{slope}=r\times\frac{\text{standard deviation of }y}{\text{standard deviation of }x}\,,$ (5)
where slope is the slope estimate of the fitted regression line and $r$ is the pearson correlation coefficient.
The posterior distribution for the slope $\beta$ is centered at 0.607 (95% CI [0.60, 0.61]), with the model’s intercept being -0.08. Hence, the expected slope, and consequently, the correlation, is positive. This suggests that as elite alignment increases, mass alignment also increases. It’s evident from the regression results that the elite tend to be more politically aligned than the general public, which aligns with common expectations. However, intriguingly, the model also offers an estimate for the “baseline alignment”, i.e. the degree of alignment among the elite even when the mass shows no alignment whatsoever. According to our model, the point estimate for the baseline alignment is approximately 0.13.
Figure 6: Each data point represents the observed alignment between pair of topics separately for the elite (x-axis) and public (y-axis). To assess and compare the strength of the relationship between these variables, bayesian linear models were fitted. The subplot illustrates the posterior distribution of the slope parameter. The linear correlation is evidently positive between the variables, and for most topics in 2023, the mass exhibits higher issue alignment than the expected posterior mean.
4 Conclusion
4.1 Discussion
@@
This study investigated whether polarized groups and their respective hierarchies have equal influence on polarization observed in online systems. We proposed a method to identify these structures within a single social system and analyzed their different impacts on two aspects of polarization: structural polarization and issue alignment-based polarization. We illustrated the use of the method by applying it to Finnish polarized networks.
@@
Our findings revealed that polarized groups had an unequal impact on overall structural polarization across all the networks studied, indicating lack of balanced systems where both opposing groups would contribute equally to the polarization score. Specifically, we found that left-leaning groups played a more dominant role in three out of five topics (climate, social, and economy) in 2019. However, by 2023, the right-leaning group became the dominant force in both climate and social politics, mostly driven by a more organized elite and greater mass amplification. Finland elected a left-leaning government in 2019 and right-leaning government in 2023 [GS23].
The hierarchy of the user in a polarized networks matters. Not only did we find that the elite consistently aligned more on distinct political topics than the mass within our networks, but we also observed that the elite had a relatively higher impact on structural polarization, as evidenced by their respective marginal values. Additionally, we discovered that all networks, which experienced an overall increase in polarization during the study period, had at least one elite group that increased its cohesion. This reminds us that a network’s increased polarization does not necessarily depend on greater public amplification. Rather, it suffices for the elite to become more organized, as exemplified for the right-leaning group in the education network.
Elite are significantly more aligned than the public online, and the alignment intensified over the four years across all the topics. The absolute difference in the average level of elite alignment more than doubled over the same time span. We asked whether the public online is somehow immune to the increasingly polarized environment. Although this study cannot respond, whether the drastic increase in the mass alignment is due to the elite being more polarized, our results clearly suggests that certain ideological realignment has taken place among the public too. The relative increase in the issue alignment was more pronaunced for the mass (+228 %) than for the elite (+121 %). We also observed that in 2023, a significant number of observations clustered above the mean for various topics, suggesting a stronger association between elite and mass alignment. Hence, there is evidence that the mass has at least not remained isolated from the trend towards increased alignment.
Again, while our analysis does not definitively pinpoint whether the elite drove this increase in mass alignment, it’s clear that socially influential figures are becoming more politically aligned. This trend could potentially influence the public, aligning with past research highlighting how they often take cues from elite figures. Furthermore, our model suggested that theoretically, the alignment among the elite would persist even if the mass showed no indication of aligning on issues. This could suggest that maintaining predictable stances on a range of political topics serves as a distinguishing feature (or strategy) that sets the elite apart from the general public.
Our contribution does not address the concept of an “ideal polarized network”. Would such a network be non-hierarchical, or would it exhibit balanced contributions from both opposing parties? It is important to acknowledge that some degree of disagreement is not only inevitable but also essential in a healthy democracy. Ideally, disagreement fosters debate, encourages diverse perspectives, and promotes critical thinking, which can be deemed as fundamental components of democratic processes. How these potentially desirable dynamics would be reflected in the network structure is an open question. However, we believe it is crucial to recognize the prevalent imbalances we have in today’s polarized networks and to take them into account in the polarization-related literature. Our study also reveals a concerning trend: a rapid decline in communication patterns and a sharp rise in issue alignment, both which have happened within a relatively short timeframe.
Social media is sometimes seen as a decentralized platform for content creation [TLT+18, LPBF19] that is not monopolized by a few influential entities. However, our observations could challenge this perception. The elite groups consisted of very small number of individuals, and these were literally in the core of the polarized networks, making them structurally dominant. Yet they had greater marginal impact on polarization and were significantly more politically aligned. This raises questions about the true decentralization of social media platforms and whether they genuinely represent the diversity of political voices in our societies.
4.2 Limitations
While our study provides valuable insights into some structural mechanisms of polarization, it is not without limitations.
Many influential figures use Twitter, and it is popular social media platform among the public too. However, while the core-periphery concept has been extensively employed to model hierarchies in networks, and our manual check indicated that a majority of political candidates fell within the identified cores, it’s essential to recognize that these findings may not necessarily extend to the behavior of the “actual” political elite (which is also relatively vague concept) in society. The same applies to the mass, which represented the non-elite nodes in the network. However, it’s worth noting that, as highlighted in previous studies, Twitter users are often perceived as more politically engaged than an average citizen. This limitation is interesting, as it can imply that the mass (and the elite) outside Twitter might be even less aligned than the elite on social media, or the other way around. The function of the public within a polarized network was simplified to act as amplifiers of the messages produced by the cores. While this simplification is not uncommon in similar studies, we believe that the temporal interactions and dynamics between a network’s cores and peripheries are more complex.
While each method presented here could be easily extended to characterize polarization in the case of multiple groups, we have chosen to concentrate on the typical settings involving two groups to facilitate comparisons between the networks.
Considering network ties as directed could enhance the accuracy of network decomposition, as the direction of endorsements is not random. Assigning greater importance to nodes with an in-degree larger than their out-degree may be in some cases desirable. Likewise, when an elite endorses someone who does not reciprocate, it might suggest a special user, such as an organizational account. In this work, we interpretted the edge between two nodes more as a mutual agreement than directed endorsement.
4.3 Future work
We mainly focused on studying structrual differences between the polarized groups and embedded hierarchies in polarized networks. Building upon the insights gained from this study, future research could investigate deeper the content and attitudes within cores, and how they reflect to the structural properties of the network. Political discussions are often perceived as a clash between the status quo and change. Do these respective groups exhibit similar patterns across all networks? Can the structural properties alone of a polarized group offer insights into the positions held by its members? How are these connected to the forces driving polarization online?
Social media platforms have experienced a surge in polarization in recent years, with conflicting findings on whether it is the public or the elite driving this phenomenon. We do not claim that the methods presented here provide a direct answer to this question. However, by conceptualizing the components linked to structural polarization, this study facilitates the examination of temporal changes. Particularly when there are data available from a longer period, it becomes possible to detect persistent structures and identify strengthening groups, whether in terms of mass amplification or elite cohesion. These factors could contribute to a more nuanced understanding of the possible under- or overrepresentation of specific groups. In fact, the introduced polarization spectrums enable us to compare how different social media platforms exhibit disparities in power structures and dynamics among competing political groups. Furthermore, treating elite and mass polarization as distinct processes could assist researchers in detecting suspicious activities that contribute to polarization in online networks—such as abrupt spikes in mass amplification that deviate from their expected trajectory.
@@ 4.3 Future work
Appendix A Appendix
A.1 Political actors in cores versus peripheries
@@
A.2 Public Amplification
Figure 7: To view the core-periphery interactions as public amplification, we confirm that the majority of connections are directed towards the core. In all networks examined, most of the links between the core and periphery consist of retweets originating from the outer periphery. This proportion remains relatively consistent across different polarized groups. Therefore, it’s reasonable to characterize this dynamic as public amplification.
A.3 Political parties presence in inferred polarized groups
@@
A.4 Characterics of polarized groups
Group A Group B
Topic C P E Coh. Ampl. P Coh. C P E Coh. Ampl. P Coh. B
CLI ’19 2.2 42.4 16.6 40.1 8.6 2.8 53.0 4.0 16.8 5.2 8.7
CLI ’23 3.6 65.8 6.3 12.9 4.7 1.7 29.0 11.9 52.0 9.1 3.1
IMM ’19 3.7 71.8 2.2 7.6 2.2 2.5 22.1 29.9 50.9 5.2 1.9
IMM ’23 3.7 51.8 9.9 17.6 4.0 3.6 41.0 20.5 39.1 5.7 3.1
SOC ’19 1.5 39.0 9.9 37.1 9.8 4.4 55.1 7.2 18.2 6.7 11.3
SOC ’23 4.2 57.3 8.4 15.6 6.7 2.7 35.8 13.7 42.0 8.7 4.9
ECO ’19 1.1 16.2 8.2 54.2 17.8 4.7 78.0 2.2 10.2 4.3 3.1
ECO ’23 2.4 28.9 23.8 36.0 8.5 4.1 64.7 7.3 17.2 4.1 3.2
EDU ’19 4.9 81.0 2.9 9.1 3.9 1.2 13.0 10.9 49.1 19.8 4.4
EDU ’23 4.0 60.4 5.6 14.2 6.0 2.4 33.2 16.8 45.1 8.3 4.0
Table 4: All values in the table represent percentages. ’C’ indicates the size of the corresponding core, while ’P’ indicates the size of the corresponding periphery, divided by the total number of nodes in the network. For the polarization spectrum components, ’E/P Coh.’ (Cohesion) column reflects either elite or public cohesion, while the ’Ampl.’ (Amplification) column shows the public amplification values. Bridge denotes the component representing the overall inter-communication between groups A and B. See more details in the caption for Figure 3.
A.5 Power dynamics
@@
Figure 9: Evolution of power structures for social and education networks. See caption of Figure 4 for more details.
A.6 Marginal polarization
How much does the overall interactional polarization increase if you added an “average” core (or periphery) node to group A? Assume that an average core node has $k_{c_{A}}$ links to other core nodes in the same group, it receives $k_{cp_{A}}$ links from the periphery and forms $k_{out}$ links to the other group B. The marginal polarization could be written then as
$\Delta P_{AEI}(\text{+ corenode to A})$ $=\frac{\frac{(I_{c_{A}}+k_{c_{A}}+I_{cp_{A}}+k_{cp_{A}}+I_{p_{A}})}{\binom{n_{A}+1}{2}}+\frac{(I_{c_{B}}+I_{cp_{B}}+I_{p_{B}})}{\binom{n_{B}}{2}}-2\times\frac{E+k_{out_{B}}}{(n_{A}+1)n_{B}}}{\alpha^{\prime}}$
$\quad-\frac{\frac{(I_{c_{A}}+I_{cp_{A}}+I_{p_{A}})}{\binom{n_{A}}{2}}+\frac{(I_{c_{B}}+I_{cp_{B}}+I_{p_{B}})}{\binom{n_{B}}{2}}-2\times\frac{E}{n_{A}n_{B}}}{\alpha}$ (6)
where $I_{\bullet}$ is the observed number of links within the corresponding structure, and $E$ denotes the observed number of links between groups $A$ and $B$. Each term is adjusted for density, as described in the main text. We approximate $\alpha^{\prime}=\alpha$. Hence we get for average core node the following marginal:
$\Delta P_{AEI}(\text{+ corenode to A})$ $=\frac{\frac{(I_{c_{A}}+k_{c_{A}}+I_{cp_{A}}+k_{cp_{A}}+I_{p_{A}})}{\frac{1}{2}n_{A}(n_{A}+1))}-2\times\frac{E+k_{out_{B}}}{(n_{A}+1)n_{B}}-\frac{(I_{c_{A}}+I_{cp_{A}}+I_{p_{A}})}{\frac{1}{2}n_{A}(n_{A}-1))}+2\times\frac{E}{n_{A}n_{B}}}{\alpha}$ (7)
$\approx\frac{\frac{(I_{c_{A}}+k_{c_{A}}+I_{cp_{A}}+k_{cp_{A}}+I_{p_{A}})}{\frac{1}{2}n_{A}(n_{A}))}-2\times\frac{E+k_{out_{B}}}{n_{A}n_{B}}-\frac{(I_{c_{A}}+I_{cp_{A}}+I_{p_{A}})}{\frac{1}{2}n_{A}(n_{A}))}+2\times\frac{E}{n_{A}n_{B}}}{\alpha}$ (8)
$=\frac{\frac{(k_{c_{A}}+k_{cp_{A}})}{\frac{1}{2}n_{A}(n_{A}))}-2\times\frac{k_{out_{B}}}{n_{A}n_{B}}}{\alpha}$ (9)
And for the others:
$\Delta P_{AEI}(\text{+ core node to B})$ $\approx 2\times\frac{\frac{(k_{c_{B}}+k_{cp_{B}})}{(n_{B})^{2}}-\frac{k_{out_{A}}}{n_{B}n_{A}}}{\alpha}$ (11)
$\Delta P_{AEI}(\text{+ periphery node to A})$ $\approx 2\times\frac{\frac{(k_{p_{A}}+k_{cp_{A}})}{(n_{A})^{2}}-\frac{k_{out_{B}}}{n_{A}n_{B}}}{\alpha}$ (12)
$\Delta P_{AEI}(\text{+ periphery node to B})$ $\approx 2\times\frac{\frac{(k_{p_{B}}+k_{cp_{B}})}{(n_{B})^{2}}-\frac{k_{out_{A}}}{n_{B}n_{A}}}{\alpha}$ (13)
A.7 Statistical testing
Our analysis requires the network to be decomposed into $k$ polarized groups (assortativity assumption) after which each polarized group is decomposed further into separate hierarchical groups (hierarchy assumption). Therefore, in case of $k=2$, the overall number of different groups would be four. From now, we assume $k=2$, and label the identified polarized groups as $A$ and $B$. For the hierarchical decomposition, we label the cores of $A$ as $c_{A}$ and its periphery as $p_{A}$ (similary for B).
We know that most existing methods for determining the modular structure of networks face issues. They often overlook the statistical evidence supporting the identified patterns, making it difficult to distinguish genuine structures from noise. In our case, we want to be sure about two things: 1) Is there enough evidence that the network can be decomposed into two groups? and 2) If so, is there enough evidence that these groups can be partitioned further into separate hierarchies?
Each network undergoes the following partitioning pipeline: 1) Network is partitioned into two blocks with the non-uniform planted partition model. 2) To find the embedded hierarchies, we consider the groups independently, and run the core-periphery algorithm provided here. We apply the traditional hub-and-spoke description of core-periphery structure as it has been shown to explain online amplification better than the layered one. We run this two-step decomposition for 100 times and save each obtained partition for the statistical significance testing.
Fist, defining a hierarchy is harder than defining the absence of it. Perhaps the most common baseline model in network science, Erdős–Rényi model, generates random graphs, where each edge $(i,j)$ is included in the graph with probability $p$ independently from every other edge. Hence, the possible heterogeneity that emerge in such graph is purely due to random fluctuations, making it decent candidate for null model.
A stricter null model would be the configuration model, another common option for testing the statistical evidence of an observed structure. Essentially, it is a mathematical model used to generate random networks with a specified degree sequence. Interestingly, previous research has shown how the degree sequence of a network can fully explain the core-periphery structure in certain situations [KM18]. We selected ER-model over the configuration model for three reasons:
  1. 1.
    We use the core-periphery structure to identify the structurally dominant users, and it is acceptable for this significance to stem from a user’s higher degree.
  2. 2.
    Some work (e.g. [YS23]) has discussed the possible shortcomings of using configuration model in detecting core-periphery structures, claiming it likely leading to low power for the statistical test. They illustrate this with an ideal core-periphery structure, for which the same original network is the only one that maintains the fixed degree sequence. Therefore, no rewiring is possible, resulting in a hypothesis test yielding a p-value of 1. [KM18] claims that an additional block, such as a community or another core-periphery, is required to conclude the significance of the observed core-periphery. However, as far as we know, there isn’t a generative model capable of inferring multiple core-periphery structures simultaneously, and such model is typically required to measure the description length.
  3. 3.
Ideally, we would apply directly a constrained version of nested stochastic block model [Pei14a] because of the hierarchical nature of our overall decomposition approach. Unfortunately, it is out of scope of this work to develop such method, and hence we employ the method described in the main text for model selection.

### v2

@@ 4.4 Elites are Significantly More Aligned than the Masses
Elite members are consistently more aligned than the public in all topics studied here, as shown in Fig. 4A. Higher alignment suggests that the elite regularly maintain more predictable stances on various political themes. In 2019, among both elites and masses, topic pairs immigration-education and immigration-climate were the most aligned. Conversely, the lowest alignment was found between the topics immigration-social security and education-social security.
@@
After four years the gap between elite and mass alignment widened—not only did the elite show a much higher level of issue alignment (the mean level doubled), but the disparity between the masses and elites also increased in 2023. The overall level of mass alignment did not stay at the same level as in 2019 either. Instead, the mass also experienced a notable increase in alignment across all topics, as the mean line, corresponding to the average of alignment over all the topics, rose from 0.13 to 0.44. For the elite groups, eight out of ten pairs of topics had an NMI value over 0.8, which can be considered very high. Results show that knowing elite member’s stance on climate politics would reveal the most information on their views about immigration policies compared to the other topics included in this study. Interestingly, climate-economy experienced the smallest increase in alignment.
4.5 Elites and Masses of Opposing Groups Show Distinct Activity Patterns
Figure 5: Activity patterns in the most polarized networks in 2023 separated at group and hierarchy level. In all networks, largest jump in activity takes place approximately three weeks before the election day, particularly within right-leaning group in immigration network. Which opposing group is more active depends on the issue. For instance, activity within the right-leaning elites is substantially higher in immigration and economy, whereas in climate, social security and education, left-leaning elites appear to be more active. The extent of activity of a specific group does not translate into the observed polarization. Figures for remaining topics and for 2019 can be found in Appendix G.
So far, we have focused on aggregate networks, meaning we constructed networks based on all interactions that occurred during the specified time period. These networks were unweighted, as there are no established guidelines for computing polarization in weighted graphs (or whether it should be done at all).
A polarized group might exhibit high cohesion among elites or the masses, yet show non-uniform endorsement patterns, suggesting that once-formed ties are not consistently maintained. Similarly, peaks in mass amplification may occur as bursts, while otherwise following stable patterns. To assess the temporal dynamics of interactions between elites and masses, we labeled each interaction according to the identified groups and their hierarchies (see Section 3.3), and plotted them leading up to the election day (see Fig. 5).
The most crucial finding is that a higher level of activity does not dictate group’s structural contribution to the observed polarization. For instance, in the climate network, the elite cohesion of the right-leaning group was approximately twice as large as that of the left-leaning elite in 2023. Yet, left-leaning elites consistently endorsed each other more than right-leaning elites throughout the entire time period. However, in immigration network, where right-leaning group also dominated in terms of elite cohesion and mass amplification in 2023, also showed more active elites and masses.
For the most polarized networks in 2023 (climate, immigration, and economy), it is apparent that the activity of elite groups varies depending on the issue. In some topics, such as immigration and economy, mass-elite and elite-elite activity are closely matched for both groups. In others, like climate, mass-elite interactions can substantially exceed elite-elite interactions, particularly within the right-leaning group.
5 Discussions and Conclusions
@@
This study examined the differential influence of polarized groups and their hierarchies on online polarization. We developed a method to identify these structures within a single social system and analyzed their impacts on structural polarization and issue alignment-based polarization. Applying this method to Twitter networks during Finnish parliamentary elections yielded revealing insights.
@@
Our findings demonstrated an unequal impact of polarized groups on overall structural polarization across all studied networks, challenging the notion of balanced systems where opposing groups contribute equally to polarization scores. In 2019, left-leaning groups dominated in three out of five topics (climate, social security, and economy). However, by 2023, right-leaning groups became the dominant force in climate and social security networks, primarily due to a more organized elite and greater mass amplification. These shifts align with Finland’s political transitions: a left-leaning government in 2019 and a right-leaning one in 2023 [25].
User hierarchy in polarized networks is crucial. Our findings reveal that elites consistently demonstrate higher alignment on distinct political topics and exert relatively greater influence on structural polarization compared to the masses. Networks experiencing increased overall polarization invariably showed at least one elite group with enhanced cohesion, underscoring that heightened polarization doesn’t necessarily require greater mass amplification. The right-leaning group in the education network exemplifies how elite organization alone can increase polarization. We also observed how elite and mass alignment more than doubled over the four years. The relative increase in issue alignment was more pronounced for the masses (+228%) than for the elites (+121%), suggesting that the public has not remained insulated from the trend towards increased alignment.
We also examined activity patterns among elites and masses, as well as between opposing groups. Our findings revealed that high activity levels in certain subgroups do not necessarily correlate with increased observed polarization. While extending current polarization scores to account for user activity is possible, the optimal method for doing so remains unclear in polarization literature. Consequently, we analyzed group and hierarchical activities separately, proving the complexities we have in quantifying political polarization.
Our work helps make sense of today’s polarized environments. The overrepresentation of elite voices in traditional and social media may skew public perception of polarization, potentially exacerbating feelings of division. We provide approaches to precisely quantify the degree of this phenomenon by comparing elite polarization to that of the surrounding public in online spaces. Furthermore, by treating elite and mass polarization as distinct processes, our method enables the detection of anomalous patterns, such as unexpected surges in mass amplification.
@@
Appendix A Marginal Polarization
How much does the overall structural polarization increase if you added an “average” core (or periphery) node to group A? Assume that an average core node has $k_{c_{A}}$ links to other core nodes in the same group, it receives $k_{cp_{A}}$ links from the periphery and forms $k_{out}$ links to the other group B. The marginal polarization could be written then as
$\Delta_{c_{A}}P_{AEI}$ $=\frac{\frac{(I_{c_{A}}+k_{c_{A}}+I_{cp_{A}}+k_{cp_{A}}+I_{p_{A}})}{\binom{n_{A}+1}{2}}+\frac{(I_{c_{B}}+I_{cp_{B}}+I_{p_{B}})}{\binom{n_{B}}{2}}}{\alpha^{\prime}}$
$\quad-2\times\frac{E+k_{out_{B}}}{(n_{A}+1)n_{B}}$
$\quad-\frac{\frac{(I_{c_{A}}+I_{cp_{A}}+I_{p_{A}})}{\binom{n_{A}}{2}}+\frac{(I_{c_{B}}+I_{cp_{B}}+I_{p_{B}})}{\binom{n_{B}}{2}}}{\alpha}$
$\quad-2\times\frac{E}{n_{A}n_{B}}\,,$ (5)
where $I_{\bullet}$ is the observed number of links within the corresponding structure, and $E$ denotes the observed number of links between groups $A$ and $B$. Each term is adjusted for density, as described in the main text. We approximate $\alpha^{\prime}=\alpha$. Hence, we get for average core node the following marginal:
$\Delta_{c_{A}}P_{AEI}$ $=\frac{\frac{(I_{c_{A}}+k_{c_{A}}+I_{cp_{A}}+k_{cp_{A}}+I_{p_{A}})}{\frac{1}{2}n_{A}(n_{A}+1)}-2\times\frac{E+k_{out_{B}}}{(n_{A}+1)n_{B}}}{\alpha}$
$\quad-\frac{\frac{(I_{c_{A}}+I_{cp_{A}}+I_{p_{A}})}{\frac{1}{2}n_{A}(n_{A}-1)}+2\times\frac{E}{n_{A}n_{B}}}{\alpha}$
$\approx\frac{\frac{(I_{c_{A}}+k_{c_{A}}+I_{cp_{A}}+k_{cp_{A}}+I_{p_{A}})}{\frac{1}{2}n_{A}(n_{A})}-2\times\frac{E+k_{out_{B}}}{n_{A}n_{B}}}{\alpha}$
$\quad-\frac{\frac{(I_{c_{A}}+I_{cp_{A}}+I_{p_{A}})}{\frac{1}{2}n_{A}(n_{A})}+2\times\frac{E}{n_{A}n_{B}}}{\alpha}$
$=\frac{\frac{(k_{c_{A}}+k_{cp_{A}})}{\frac{1}{2}n_{A}(n_{A})}-2\times\frac{k_{out_{B}}}{n_{A}n_{B}}}{\alpha}$
Similarly, for the remaining node types:
$\Delta_{c_{B}}P_{AEI}$ $\approx\frac{2}{\alpha}\times(\frac{(k_{c_{B}}+k_{cp_{B}})}{n_{B}^{2}}-\frac{k_{out_{A}}}{n_{B}n_{A}})$ (7)
$\Delta_{p_{A}}P_{AEI}$ $\approx\frac{2}{\alpha}\times(\frac{(k_{p_{A}}+k_{cp_{A}})}{n_{A}^{2}}-\frac{k_{out_{B}}}{n_{A}n_{B}})$ (8)
$\Delta_{p_{B}}P_{AEI}$ $\approx\frac{2}{\alpha}\times(\frac{(k_{p_{B}}+k_{cp_{B}})}{n_{B}^{2}}-\frac{k_{out_{A}}}{n_{B}n_{A}})$ (9)
Appendix B Statistical Testing
Our method requires the network to be partitioned into $k$ polarized groups (assortativity assumption) after which each polarized group is decomposed further into separate hierarchical groups (hierarchy assumption). Therefore, in case of $k=2$, the overall number of different groups would be four. From now, we assume $k=2$, and label the identified polarized groups as $A$ and $B$. For the hierarchical decomposition, we label the cores of $A$ as $c_{A}$ and its periphery as $p_{A}$ (similary for B).
We know that most existing methods for determining the modular structure of networks face issues. They often overlook the statistical evidence supporting the identified patterns, making it difficult to distinguish genuine structures from noise. In our case, we want to be sure about two things: 1) Is there enough evidence that the network can be decomposed into two groups? & 2) If so, is there enough evidence that these groups can be partitioned further into separate hierarchies?
Each network undergoes the following partitioning pipeline: 1) Network is partitioned into two blocks with the non-uniform planted partition model. 2) To find the embedded hierarchies, we consider the groups independently, and run the core-periphery algorithm provided in [21]. We apply the traditional hub-and-spoke description of core-periphery structure as it has been shown to explain online amplification better than the layered one. We run this two-step decomposition for 100 times and save each obtained partition for the statistical significance testing.
First, defining a hierarchy is harder than defining the absence of it. Perhaps the most common baseline model in network science, Erdős–Rényi model, generates random graphs, where each edge $(i,j)$ is included in the graph with probability $p$ independently from every other edge. Hence, the possible heterogeneity that emerge in such graph is purely due to random fluctuations, making it decent candidate for null model.
A stricter null model would be the configuration model, another common option for testing the statistical evidence of an observed structure. Essentially, it is a mathematical model used to generate random networks with a specified degree sequence. Interestingly, previous research has shown how the degree sequence of a network can fully explain the core-periphery structure in certain situations [33]. We selected ER-model over the configuration model for three reasons:
  1. 1.
    We use the core-periphery structure to identify the structurally dominant users, and it is acceptable for this significance to stem from a user’s higher degree (as in e.g. [22]).
  2. 2.
    Some work (e.g. [65]) has discussed the possible shortcomings of using configuration model in detecting core-periphery structures, claiming it likely leading to low power for the statistical test. They illustrate this with an ideal core-periphery structure, for which the same original network is the only one that maintains the fixed degree sequence. Therefore, no rewiring is possible, resulting in a hypothesis test yielding a p-value of 1. [33] claims that an additional block, such as a community or another core-periphery, is required to conclude the significance of the observed core-periphery. However, as far as we know, there isn’t a generative model capable of inferring multiple core-periphery structures simultaneously, and such model is typically required to measure the description length.
  3. 3.
Ideally, we would apply directly a constrained version of nested stochastic block model [50] because of the hierarchical nature of our overall decomposition approach. Unfortunately, it is out of scope of this work to develop such method, and hence we employ the method described in the main text for model selection.
Appendix C Summary of Networks
Table 1: Retweet networks were constructed for both election years, covering five distinct networks each. These networks were constructed based on Twitter data obtained over a span of 12 weeks leading up to the respective election day, which were 14.4. for 2019 and 2.4. for 2023. $|N|$ represents the count of unique nodes (users), and $|E|$ denotes the count of unique edges (retweets) in the network after the preprocessing.
Topic $|N|_{19}$ $|E|_{19}$ $|N|_{23}$ $|E|_{23}$
Climate politics 16 639 50 605 11 264 34 509
Immigration politics 6 739 20 923 9 519 36 889
Social security 13 926 39 778 11 395 34 173
Economic policy 5 880 10 962 10 031 37 111
Education policy 9 843 20 314 12 814 35 619
Appendix D Political Actors in Cores versus Peripheries
@@
We report the values of $\gamma$ for each network in Table 2.
@@
Appendix E Public Amplification
Figure 6: To view the core-periphery interactions as mass amplification, we confirm that the majority of connections are directed towards the core. In all networks examined, most of the links between the core and periphery consist of retweets originating from the outer periphery. This proportion remains relatively consistent across different polarized groups. Therefore, it’s reasonable to characterize this dynamic as mass amplification.
Appendix F Political Parties in Inferred Polarized Groups
We verified that the inferred groups can be called either left-leaning or right-leaning based on the real political affiliations of the political actors in the groups in Fig. 7.
@@ Appendix F Political Parties in Inferred Polarized Groups
Appendix G Activity Patterns
@@
Figure 8: Activity patterns in social security and education networks 2023. See caption of Figure 5 for more details.
Figure 9: Activity patterns in climate, immigration and economy networks 2019. See caption of Figure 5 for more details.
Figure 10: Activity patterns in social security and education networks 2019. See caption of Figure 5 for more details.